OpenStudy (yrivers36):

Factor completely 8x^3-27 please give explanation.

OpenStudy (anonymous):

(a-b)(a^2 + ab + b^2) (2x - 3)(4x^2 + 6x + 9)

OpenStudy (yrivers36):

Im confused as to how you got the 2x

OpenStudy (anonymous):

8x^3-27 = (2x)^3 - (3)^3 Now using the identity a^3 - b^3 = (a - b) (a^2 + ab + b^2) (2x)^3 - (3)^3 = (2x - 3)[(2x)^2 + (2x)(3) + (3)^2] = (2x - 3) (4x^2 + 6x + 9)

OpenStudy (anonymous):

Is everything clear now..???

OpenStudy (yrivers36):

yes thank you

OpenStudy (anonymous):

How can you identify when you need to use this equation? and what you did in the beginning was use the cubed squared function right?